### v1

# Optimizing ML Concurrent Computation and Communication with GPU DMA Engines

Abstract
Concurrent computation and communication (C3) is a pervasive paradigm in ML and other domains, making its performance optimization crucial. In this paper, we carefully characterize C3 in ML on GPUs, which are most widely deployed for ML training and inference. We observe that while C3 leads to performance uplifts, the uplifts are far lower than ideal speedups (serial computation and communication versus maximum of computation or communication; all times from isolated executions). C3 on average achieves only 21% of ideal speedup, this is due to known challenges of compute and memory interference between concurrent GPU kernels (that is, sharing of GPU’s compute units, caches and HBM).
To attain better performance for C3, first, we evaluate dual strategies of schedule prioritization and careful resource partitioning of compute units on GPUs to push performance attained with C3 (on average 42% of ideal speedup). We also provide heuristics that can guide a runtime while employing these strategies. To further enhance C3 performance, we propose to mitigate C3 interference by offloading communication tasks to the GPU’s DMA engines. To this end, we build Concurrent Communication CoLlectives (ConCCL) proof-of-concepts that harness DMA engines for communication. We show how ConCCL considerably closes the gap between realized and ideal speedup for C3 (on average 72% of ideal speedup is realized, up to 1.67 speedup). Overall, our work makes a strong case for GPU DMA engine advancements to better support C3 on GPUs.
Index Terms:
Concurrency, DMAs, GPU, MLI Introduction
Large-scale machine learning (ML) models continue to harness distributed computing over increasingly large clusters of GPUs. A consequence of this is intermingling of computation and communication which, in absence of data dependencies, can be executed concurrently on the GPU. There are several ML algorithmic choices that lead to concurrent computation and communication, termed C3 in this work, such as data-parallelism [1], fully sharded data parallel (FSDP) [2], nano-batching [3] and more. As such, characterization of and optimization for C3 on GPUs is important.
To achieve this, we first create a taxonomy of C3 and provide detailed characterization of C3 anchored on this taxonomy. Given our ML focus, we study C3 manifestations wherein the computation kernel is a matrix-matrix multiplication (or GEMM) kernel and the communication kernel is a collective operation such as an all-gather among multiple GPUs. By analyzing C3 scenarios from the training of LLaMA-70B and LLaMA-405B [4] models, along with some synthetic C3 scenarios, we provide broad coverage for our proposed taxonomy.
Next, using detailed experiments, we determine the compute and memory bandwidth requirements of isolated executions of the computation and communication kernels under study. This enables us to get an assessment of ideal performance uplift via concurrency. We observe that, when it comes to compute needs, on a state-of-art GPU such as AMD Instinct™ MI300X which has ample compute units (or GPU compute cores), communication kernels need only up to 10-20% of available units making the rest available to concurrent computation kernels (Figure 1, left). GEMMs, as expected, manifest a spectrum of behaviors, some resilient to losing compute units to concurrent communication while the rest being sensitive. On the memory front, except for memory bandwidth-bound GEMM kernels, we observe that the bandwidth needs of compute-bound GEMMs and communication kernels can be met in tandem by the high memory bandwidth made available by MI300X GPUs. Overall, for C3 scenarios under study, compared to a baseline which serializes computation and communication, if the shorter of the computation or communication kernels is completely hidden, ideal speedups of 1.6 (average) and 2 (maximum) are possible.
However, here we observe that of this ideal speedup, on average only 21% (1.13) is realized. This is not unexpected, as concurrent GPU kernels cause mutual interference with each other as they share compute cores, caches, high bandwidth memory (HBM) bandwidth and more. That is, available resources to a given GPU kernel are lower in presence of other kernels than in isolation. In fact, prior work [5] has observed that due to such interference, C3 can even lead to slowdowns and loss of ML throughput.
To bridge the above performance gap, we first evaluate two strategies: schedule prioritization and careful partitioning of GPU compute units. Specifically, using our isolated execution analysis above, we observe that prioritizing scheduling of kernels with lower resource requirements prevents starvation and leads to overall better performance. Similarly, we also demonstrate that the resource partitioning available on MI300X GPUs wherein certain compute units on GPUs can be exclusively marked for a given stream of work can be harnessed to push C3 performance. Overall, using these strategies we can push performance attained with C3 to an average of about 42% of available ideal speedup. Additionally, we also provide heuristics that can guide a runtime in employing these strategies.
Finally, to further enhance C3 performance, we tackle the interference incurred with C3 manifestations today. That is, for compute interference, instead of splitting available compute units among compute and communication kernels (Figure 1, left), we harness existing direct memory access (DMA) engines on MI300X GPU and offload communication to them making all compute units available to concurrent GEMM computations (Figure 1, right). Given the placement of DMA engines in MI300X GPU (Section II-A), this has the added advantage of lowering interference in subset of caches as well. To this end, we build ConCCL proof-of-concepts (PoCs) for ML collectives such as all-gather and all-to-all111Note that, as GPU DMA engines do not support arithmetic operations, we do not consider offloading all-reduce kernels to DMAs.. We demonstrate that our simple PoCs are at par with existing communication libraries for bandwidth-bound scenarios while also reducing interference in C3, leading to attainment of on average 72% of ideal speedup, up to a max of 1.67 speedup. Overall, our work makes a strong case for continued investment and betterment of GPU DMA engines and we conclude with a discussion of further GPU enhancements to support C3 efficiently.
The key contributions of this work are as follows:
As concurrent computation and communication (C3) is an important paradigm in ML, high performance computing (HPC) and other domains, we present a detailed taxonomy and characterization of this paradigm on GPUs.
Our analysis shows that while C3 can lead to performance uplifts compared to serial execution, not all of the potential speedup is realized (21% of ideal speedup is realized). This is not unexpected, as compute and memory interference among concurrent kernels causes slowdowns.
To address the above gap, we first observe that C3 performance uplifts can be improved (42% of ideal speedup) by dual strategies of schedule prioritization and careful resource partitioning of compute units on GPUs.
To attain further performance uplifts, unlike current state-of-art communication libraries which offload communication to GPU compute-units, we build ConCCL (Concurrent Communication CoLlectives) proof-of-concepts, which offload communication to DMA engines on GPUs.
Using ConCCL, we demonstrate that C3 interference is lowered, considerably closing the gap between realized and ideal speedup (on average 72% of ideal speedup is realized).
Overall, we make a strong case to enhance the capabilities of DMA engines on GPUs, as they can play an important role in efficiently supporting C3 in GPUs.
II Background
II-A AMD Instinct MI300X GPU and Compute Orchestration
In this paper, we study C3 using AMD Instinct MI300X GPU depicted in Figure 2. A single MI300X GPU employs advanced packaging to integrate heterogeneous chiplets. Specifically, each MI300X is comprised of eight accelerator complex dies (XCD) [6] vertically stacked over four I/O dies (IOD), two XCDs per IOD [6, 7].
The IODs are comprised of AMD Infinity Cache™, a shared memory-side last-level cache (LLC) which is 256MB in size. Also, the IODs contain the memory interface to the on-package HBM. Each MI300X has a total of eight HBM stacks, each with 24GB (total of 192GB) for a combined peak memory bandwidth of 5.3TB/s [8]. Additionally, the IODs also contain 14 DMA copy engines222AMD HSA runtime API [9] allows querying of available DMA engines on MI300X., termed SDMA (or system DMA) which are available for intra-node data transfers between GPUs with necessary address mapping support.
Each XCD is comprised of 38 active compute units (CUs), which are highly threaded and parallel GPU processor cores. Each XCD also has shared L2 cache of 4MB shared across all CUs within the XCD. Overall, in an MI300X GPU, there are 304 CUs. Also depicted in Figure 2, is compute orchestration. Computations are offloaded to GPUs as kernels, each comprising multiple workgroups which are scheduled on available CUs across all XCDs.
Large-scale ML (the focus of this work) often employs multiple GPUs in tandem. In this work, we focus on the AMD MI300X Infinity Platform comprised of a 8x MI300X node with a fully-connected topology. Each MI300X connects to seven other MI300X GPUs using AMD Infinity Fabric™ [8] bi-directional links (each uni-directional bandwidth of 64GB/s/link).
II-B Direct Memory Access (DMA) Engines in GPUs
As discussed above, each MI300X is comprised of 14 SDMA copy engines available for transfers between address-space shared GPUs. We depict the steps involved for users to use these DMA engines in Figure 3. Using either heterogeneous-computing interface for portability (HIP) [10] or heterogeneous system architecture (HSA) [9] runtime API calls, at user-level, a programmer requests a single data transfer to be done using SDMA engines. Under the hood, this causes the runtime on CPU to place a command packet ( ) in the DMA queue placed in system memory for a specific GPU (either source or destination GPU for the transfer). The DMA engine of the GPU polls this queue and pulls the command from the queue ( ) and processes it ( ). Once decoded, the DMA engine issues necessary reads/writes from HBM memory of source/destination GPU respectively to complete the transfer ( ).
II-C ML Operators and C3 in ML
As large-scale ML training and inference increasingly relies on GPU clusters, GPUs need to support both efficient computation and efficient communication. Furthermore, there are several ML algorithmic choices that lead to concurrent computation and communication. C3 entails a pair of computation kernel and communication kernel that have no data dependencies and as such as be scheduled concurrently on a given GPU.
Examples of such algorithmic choices include data-parallelism [1], FSDP [2], nano-batching [3] and more. In data-parallelism, back propagation phase of ML training comprises matrix-matrix multiplication gradient kernels (both input and weight gradients) which can be scheduled concurrently with reduction (communication) of weight gradients across GPUs for intra-layer concurrency (input gradient) or inter-layer concurrency (weight gradient of previous layer). Similarly techniques like FSDP, gather model weights for a given layer on a GPU (communication), while performing computations of previous layers. Finally, optimizations like nano-batching [3] break down a single batch of inputs into nano-batches opening up opportunity for compute and communication kernels from different nano-batches to be co-scheduled on a GPU. Overall, given the ample prevalence of C3 in ML, characterization of and optimization for C3 on GPUs is important and hence is the focus of this work.
III C3 Taxonomy
We begin with a taxonomy for C3 and subsequently anchor our characterizations and optimizations on this taxonomy.
As discussed in Section II-C, the focus of our work is C3 manifestations in ML. More specifically, while there exist a variety of operators/computations in ML, in this work we focus on two primary operators, namely, general matrix-matrix multiplications (or GEMM) kernels and communication kernels for they contribute to majority of the ML execution time in both training and inference scenarios across a spectrum of models [11]. Based on this, we depict our proposed C3 taxonomy in Figure 4. Specifically, we consider three key types of C3: G-long, C-long and GC-equal.
As their names suggest, G-long is a C3 manifestation where the computation kernel (or GEMM in this case) is more dominant (>%115) when comparing their isolated execution times. Similarly, C-long is a C3 manifestation where communication kernel is more dominant (>%115) in terms of execution time. Finally, GC-equal is a C3 manifestation where both kernels are comparable (within 15% of each other).
Even within this framework, we further observe that, the relative magnitude ( in Figure 4) of computation and communication kernels with regards to each other has an effect on expected interference incurred and as such performance attained. That is, when two kernels run concurrently on a GPU, they interfere with each other in both compute (splitting of compute units) and memory sub-system (caches, HBM). As such, we consider this part of our taxonomy and we vary the relative magnitudes of C3 scenarios under study in our work.
Finally, not all GEMM or communication kernels are the same. For GEMMs, we consider two broad categories of compute-bound and memory-bound ( in Figure 4) kernels. We define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise). For communication, we consider two types of commonly used ML multi-GPU collective kernels - all-gather and all-to-all. Similar to GEMM kernels, such ML collective kernels, depending on communication size, can be latency-bound or bandwidth-bound ( ). We classify a communication kernel with its associated size as latency-bound if the kernel latency at/before this size does not increase commensurate to size. We further discuss the specific sizes and C3 manifestations we consider to cover this taxonomy, the source (ML model) for these manifestations and rationale in Section IV-A.
IV Baseline C3 Characterization
IV-A Methodology
IV-A1 System Setup
As discussed in Section II-A, we study C3 in ML using the AMD MI300X Infinity Platform comprised of 8x MI300X node with a fully connected topology.
Further, recall that we focus on two primary ML operators, namely, GEMM kernels and for communication, ML collectives kernels such as all-gather and all-to-all. For former, we employ AMD ROCm™ [12] rocBLAS library [13] comprised of high-performance GEMM kernels. For ML collectives, we employ AMD ROCm Communication Collectives Library or RCCL [14], a library of standard collective communication routines for GPUs.
We use multiple GPU streams [15] or independent set of GPU kernels that can be co-scheduled on the GPU to concurrently launch GEMM (computation) and ML collective (communication) kernels by scheduling each type of kernel in its independent stream. Additionally, we leverage the feature available on MI300X GPU to reserve CUs for specific stream to study compute needs of kernels. We use rocprof [16] GPU kernel performance profiling tool to assess kernel execution times for this study. We run total of 15 executions, first 6 are warmups and then 9 are actual measured.
IV-A2 C3 Manifestations Under Study
| gemm-tag | gemm-size | source |
| cb1 | 8192x8192x8192 | LLaMA-70B |
| cb2 | 16384x8192x16384 | LLaMA-405B |
| cb3 | 16384x16384x8192 | LLaMA-405B |
| cb4 | 18432x8192x16384 | LLaMA-405B |
| cb5 | 106496x8192x16384 | LLaMA-405B |
| mb1 | 8192x57344x8192 | LLaMA-70B |
| mb2 | 16384x106496x8192 | LLaMA-405B |
We discuss the GEMM kernels we study in this work, their sizes and their sources in Table I. As listed, we source our GEMM sizes from training of LLaMA-70B and LLaMA-405B [4] models assuming a total of 8192 tokens in a batch. Further, for ease of reading, we tag these GEMM sizes/kernels as compute-bound (cb) or memory-bound (mb). Recall that, we define a kernel to be compute-bound if its measured op-to-byte ratio is larger than machine op-to-byte as calculated from peak compute and memory throughput of underlying processor (kernel is memory-bound otherwise).
We list the C3 manifestations we study in this work in Table II. We first provide a tag for C3 being studied using GEMM type followed by the array/data size of the parallel collective operation. This is listed in column C3 in the table. Further, in all our analysis, we separately present the results for different collective types (all-gather, all-to-all). We also list in the table the source for each C3 manifestation. As shown, of 15 unique C3 combinations we have, seven are manifested in training of LLaMA-70B and LLaMA-405B [4] models (we assume 8-way sharding and FSDP [2]) using all-gather collective. To provide good coverage for our proposed taxonomy in Section III, we add additional synthetic C3 manifestations, wherein, we keep the GEMM kernel size as observed in LLaMA models (Table I) but add more communication sizes and repeat all C3 scenarios for all-to-all collective as well.
| C3-type: G-long | |||
| C3 | source | C3 | source |
| mb1_896M | LLaMA-70B | mb2_3.25G | LLaMA-405B |
| mb1_4G | synthetic | mb1_6G | synthetic |
| cb3_512M | LLaMA-405B | cb4_512M | LLaMA-405B |
| cb5_1.63G | LLaMA-405B | cb4_1G | synthetic |
| C3-type: C-long | |||
| mb1_13G | synthetic | cb2_3.25G | LLaMA-405B |
| cb4_2.5G | synthetic | cb1_896M | LLaMA-70B |
| cb5_20G | synthetic | ||
| C3-type: GC-equal | |||
| mb2_26.5G | synthetic | cb5_13G | synthetic |
We also list the taxonomy for each C3 in Table II. Notice that, we have more of G-long scenarios than C-long scenarios than GC-equal scenarios. This is so, as shown in Table II, majority of C3 manifestations in LLaMA models today are of G-long type and we consciously wanted to limit the synthetic scenarios we added. That said, we provide all of our results grouped by C3 type to understand the performance for each of these different types.
IV-B Isolated Execution Characterization
We first profile the computation and communication kernels in isolation to ascertain their compute and memory needs and assess potential for performance when they are executed concurrently.
IV-B1 Compute Needs
Recall that, to execute computation and communication concurrently on GPUs, we launch two independent kernels and as such available compute cores or compute units (CUs) for MI300X (Section II-A) are split between these two kernels. This compute interference can slowdown each of these kernels. We study this in Figure 5 (a) for GEMM kernels and (b)/(c) for subset of all-gather and all-to-all sizes (rest of the sizes show similar behavior).
For GEMM kernel slowdown, we depict two of the available seven GEMMs under study (Table I) in Figure 5(a), as they represent the extremes in terms of slowdown. To plot slowdown, we compare a GEMM execution when the specified number of CUs are taken away from the kernel to an execution where all CUs are available to the kernel (0 on x-axis is when the kernel has all 304 CUs available to it). As depicted in the figure, memory-bound GEMMs are resilient to CU loss, even attaining speedups333We observe better cache behavior for this GEMM due to less number of concurrent threads. (highlighted with a circle), while compute-bound GEMMs can incur increasing slowdowns as more CUs are taken away from the GEMMs. As for communication, we do a similar slowdown analysis in Figure 5(b)/(c) and observe that unlike GEMMs, all-gather kernels need 32 CUs, while all-to-all kernels need 64 CUs beyond which there is no benefit to allocating more CUs to the kernel.
Overall, considering the compute needs of these kernels in tandem, it can be expected that memory-bound GEMMs can co-exist with communication kernels, because memory-bound GEMMs are resilient to loss of about 32-64 CUs needed by communication kernels. On the other hand, for the same CU loss, compute-bound kernels suffer up to 17-27% slowdowns.
IV-B2 Memory Needs
Similar to compute needs analysis above, we next consider memory needs of these kernels. As discussed in Section II-A, MI300X GPU employs a shared last-level AMD Infinity Cache which, being memory-side, subsumes all of HBM traffic. As such we study the memory bandwidth needs of kernels under study by providing their relative AMD Infinity Cache bandwidth utilization in Figure 6. Note that, we only show all-to-all kernels and skip all-gather kernels as the latter have about 14% lower bandwidth vs. all-to-all kernels.
As depicted in Figure 6, the bandwidth needs of memory-bound GEMM kernels dwarf that of all other kernels under study. That said, the compute-bound GEMM and communication kernels are in similar ballpark of bandwidth needs. Further, given that the combined bandwidth needs of compute-bound GEMM kernels and communication kernels still leaves headroom in peak bandwidth available, these kernels can share available bandwidth in concurrence.
IV-B3 Ideal speedup projection
Based on above isolated execution analysis of compute and memory needs, we project in Figure 7 the ideal speedup possible for C3 scenarios under study. For this analysis, we assume that the best speedup possible is when the smaller of the two kernels (computation or communication) happens completely in the shadow of the other. As depicted, wide variety of speedups are possible (1.1 to close to 2) and this is dictated by the relative magnitudes of these kernels (Section III). However, based on compute/memory needs analysis, due to compute/memory interference not all of this ideal performance can be attained.
IV-C C3 Characterization
We present characterization of C3 for different scenarios under study in Figure 8. In this figure we refer to baseline C3 performance as c3_base. Note that, with concurrent scheduling, we schedule GEMM kernel first (as scheduling both the kernels at the exact same time is unrealistic, we minimized the scheduling delay between the two through code optimizations) in c3_base executions.
In Figure 8, we present average speedups for different groupings of C3 scenarios (collectives, and taxonomy). We also show in the graph the ideal speedups possible. The figure shows that c3_base speedups range from no speedups to up to 1.3 speedups. On average, c3_base attains 1.13 speedup which is about 21% of ideal speedup. Note that, this is not unexpected. In fact, prior work [5] observed slowdowns with C3 due to mutual interference of concurrent kernels.
Further, all-to-all attains about 0-13% of ideal speedups while all-gather attains about 24-46%. The slight edge for all-gather is due to the fact the overall memory traffic (hence memory interference) and compute needs are lower for all-gather vs. all-to-all (Section IV-B1). To understand lower memory traffic of all-gather, in simplistic terms, in all-gather, each GPU begins with a single data buffer and the end state is that every GPU holds the complete aggregated data from all other GPUs. In contrast, with all-to-all, each GPU starts with distinct data buffer for every other GPU (lot more data) and concludes with each GPU receiving unique data from all other GPUs (effectively a transpose of data buffers).
V Optimizing C3 Performance
Next, to improve C3 performance reported above, we evaluate dual strategies of schedule prioritization and careful resource partitioning of compute units on GPUs.
V-A Schedule Prioritization
Note that, thus far, for concurrent scheduling, we chose to schedule GEMM kernel first. But based off our compute needs analysis, we observe that communication kernels have far lower CU needs as compared to GEMM kernels (Section IV-B1). As such, if GEMM kernels are scheduled first, the internal GPU scheduler can in some cases allocate majority of CUs to GEMM kernel leading to potential starvation. In contrast, if we employ, schedule prioritization, that is schedule communication kernel first, later scheduled GEMM kernel will definitely have compute resources to be allocated.
We perform this schedule prioritization, referred to Figure 8 as c3_sp and show the speedups attained. Across all C3 types and collectives, we observe that schedule prioritization improves the speedups attained for C3. Overall, instead of 0-13% of ideal speedup, for all-to-all, c3_sp attains 27-46% instead. Similarly, instead of 24-46% of ideal speedup, for all-gather, c3_sp attains 38-67% of ideal speedup. Overall, on average, c3_sp helps attain 42% of ideal speedup (up from 21% of c3_base).
V-B Resource Partitioning
Note that schedule prioritization is not the only way to prevent starvation of kernel with lower resource needs. We also observe that we can leverage the feature available on MI300X GPU to reserve compute units (CUs) for specific stream to more deterministically allocate CUs to communication kernels and attain similar benefits.
We show this in Figure 8 as c3_rp which adds resource partitioning (rp) to c3_base. We sweep all possible powers-of-two CU allocations for communication kernels (and consequently take CUs away from GEMM kernel) and plot the best performing one as c3_rp. As depicted, c3_rp delivers performance improvements over c3_base which is almost the performance delivered by c3_sp (41% of ideal speedup). Note, that, adding resource partitioning to c3_sp, shown as c3_sp_rp, did not improve performance any further.
V-C Runtime Heuristics for Improved C3 Performance
We believe that either schedule prioritization or resource partitioning can help improve C3 performance, while former is indeed simpler to implement. To this end, we provide some heuristics that can guide a runtime while employing them.
Heuristic for Schedule Prioritization: As runtimes launch GPU kernels, they can use the information about number of workgroups per kernel (Section II-A) as a proxy for CU requirement of a kernel. Using this information, excepting any other prioritization that supersedes, the runtime can employ scheduling order in the order of resource requirements (number of workgroups), low to high.
Heuristic for Resource Partitioning: For resource partitioning across concurrent kernels, we describe a runtime heuristic to allocate CUs across concurrent kernels. To do so, we picked one memory-bound GEMM kernel, one compute-bound GEMM kernel, and two all-gather and all-to-all collective sizes (latency-bound, bandwidth-bound) and used the compute need analysis done in Section IV-B1 to build a lookup table of potential slowdowns when the kernel loses given number of CUs. Note that, for a given GPU this is to be done once.
Next, for any C3 scenario, we scale roofline GEMM and communication times by these slowdowns for different number of CUs to identify the CU allocation that leads to best performance, that is max(GEMM, communication) time is lowest. For roofline times, we simply focus on peak compute, memory and network throughputs and assume 70% efficiency (taking the average of our/observed [17] peak compute, memory and network efficiencies). We observe that this simple heuristic, predicts CU allocation necessary for 24 of 30 C3 scenarios. For the rest, in comparison to sweeping all possible CU allocations, our heuristic, at best loses 1.5%.
VI Optimizing C3 with ConCCL
VI-A Motivation for DMA Offloads of Communication
For co-scheduled computation and communication, the performance limiters are the compute and memory interference incurred by sharing of compute and memory resources between concurrent kernels. That is, a GEMM kernel in isolation could have been allocated all of 304 compute units (CUs) available on MI300X. In C3, however, number of CUs allocated will be lower causing what we term as compute interference and leading to lower speedups. Similar interference also occurs in the memory sub-system (e.g., caches, HBM).
One way to tackle compute interference completely and memory interference partially, is to offload communication to DMA engines on MI300X. By doing so, we can free up all available compute units for concurrent computation kernel. Further, as DMA engines are placed in IOD beyond L2 caches (Section II-B), this also eliminates any L1/L2 portion of memory interference. These dual benefits motivate us to offload communication to DMA engines on MI300X.
VI-B ConCCL: DMA-collectives PoCs
With the above strong motivation, we build ConCCL proof-of-concepts (PoCs), which are DMA-based collectives wherein we offload the collectives to DMA engines on MI300X. As MI300X DMA engines do not expose any computational functionality, we only build these PoCs for all-gather and all-to-all (and not all-reduce collective which involves reduction).
We harness the fully-connected topology of MI300X to keep the design of collectives simple. That is, we break down the collective operation into a series of individual transfers (going to different GPUs) and we schedule each such transfer on a specific available DMA engine on MI300X (recall that 14 engines are available). We leverage AMD HSA API call hsa_amd_memory_async_copy_on_engine [18] to schedule a single transfer at a given engine. Also, we implement a simple direct algorithm for our collectives. As an example, for all-gather this means that our collective implementation comprises a single step: every GPU reads the data buffer it owns and writes it to every other GPU. Additionally, unlike RCCL communication kernels which are orchestrated by launching GPU kernels, as discussed in Section II-B, DMA engine transfers are orchestrated by CPU.
Finally, we term ConCCL as PoCs for our goal in this work is not to build a high-performant communication collectives library but to use these PoCs to analyze if offloading concurrent communication collectives to DMA engines helps improve speedups possible with C3. We leave building a high-performance DMA-based collectives library to future work.
VI-C Isolated ConCCL Characterization
We first provide a comparison of our PoCs to state-of-art RCCL communication library on MI300x. We show this comparison in Figure 9. We follow the same setup details as outlined in Section IV-A1 and model multiple warm-up executions before actual measured executions.
As shown in Figure 9, our simple PoCs are slower than RCCL library for <32MB collective size by as much as 4. This is so, as it can be costly to both launch transfers from CPU to DMA engines on GPU and synchronize with CPU once the transfers are done. This launch/sync cost is not amortized for smaller sizes. That said, for larger sizes, our proposed all-gather PoCs is at par with RCCL library. Finally, recall that, as the smallest communication size we consider in our C3 scenarios is 128MB (Table II), evaluating C3 with RCCL or ConCCL is a fair evaluation as in this region the performance of both is at-par.
VI-D C3 with ConCCL Characterization
Next, we evaluate the speedups of C3 with ConCCL over sequential execution. We depict this in Figure 10.
We first compare c3_base configuration, which is C3 without schedule prioritization or resource partitioning, to ConCCL without such optimizations. As shown, across the board, ConCCL delivers higher speedups in comparison to c3_base. This is attributed to lowering the compute interference incurred by GEMM kernels in c3_base. Additionally, GEMM kernels also do not incur any memory interference in any higher level of caches such as L1/L2 with DMA offloads.
Overall, while c3_base attains 21% of available ideal speedup, with DMA offloads, ConCCL attains 66% of ideal speedup. Further, ConCCL benefits are even more pronounced for all-to-all (c3_base: 1.05, ConCCL: 1.43) which has higher compute needs (64 CUs) and as such incurs more compute interference and also has higher memory traffic/interference in higher level caches.
VI-E Schedule prioritization with ConCCL
Unlike c3_base, where both computation and communication are scheduled on GPU CUs making schedule prioritization important, with ConCCL, as communication is scheduled to DMA engines and computation to CUs (two separate entities), schedule prioritization is not necessary for ConCCL.
VI-F Resource partitioning with ConCCL
Finally, we consider resource partitioning for ConCCL. Since ConCCL does not use any compute units for communication, the allocation decision here is different from c3_base. Here we observe that, only for memory-bound GEMM kernels, as depicted in Figure 5, taking CUs away from GEMM kernel can lead to performance improvement due to improved cache behavior. Such improvement can also aid in C3 runs where communication is offloaded to DMA engines.
To study this, only for memory-bound GEMM kernels, we create a variant of ConCCL, ConCCL_rp and depict its performance in Figure 10. This variant performs slightly better than ConCCL. Point to note that, both ConCCL and ConCCL_rp performs significantly better than c3_best variant (best of all c3 variants in Section IV). We see here that on average, while c3_best attains 48% of ideal speedup, ConCCL attains 66% and ConCCL_rp attains 72% of ideal speedup, considerably closing the gap between attained and ideal speedup.
VI-G Runtime Heuristic for Resource partitioning with ConCCL
For resource partitioning with ConCCL, we recommend a much smaller subset of steps as proposed in Section V-C for baseline C3. Specifically, only the CU-loss slowdown table for a memory-bound GEMM kernel ought to be created. This table helps the runtime identify the optimal number of CUs to remove from concurrent memory-bound kernels for maximum speedups. In our analysis, for MI300X, taking away eight CUs lead to speedups for memory-bound GEMM kernels.
VII Discussion
VII-A System Evolution for Efficient C3
While ConCCL leads to more efficient C3 on GPUs, we believe additional system evolution can further help in this regard. We highlight some potential techniques below.
VII-A1 Addressing Memory Interference in C3
Our approach mitigates compute interference by offloading communication to the GPU’s DMA engines. On the MI300X GPU, DMA transfers inherently do not interact with the L1 and L2 caches, effectively eliminating cache interference at these levels. That said, contention for AMD Infinity Cache and HBM bandwidth remains, impacting performance. We leave to future work exploration of techniques such as memory-bandwidth partitioning via memory space (channel) partitioning amongst kernels, memory-aware scheduling [19] to explicitly tackle memory interference. Specifically for latter, C3 taxonomy hints to hardware can aid in memory traffic prioritization.
VII-A2 Extending ConCCL to All-Reduce Collective
Our ConCCL PoCs focus on offloading all-gather and all-to-all collectives to DMA engines. However, all-reduce operations involve both communication and computation (e.g., summing values across GPUs) and since DMA engines do not currently support arithmetic operations we could not offload all-reduce collectives to DMA engines. That said, all-reduce is also used in ML and offloading this to DMA engines can be useful. Investigating addition of arithmetic units to DMA engines is a possibility although the area/power costs of doing so ought to be balanced with possible returns. Alternately, a hybrid approach can be followed. That is, as all-reduce is comprised of a reduce-scatter and all-gather operation, only for latter DMA engines can be harnessed.
VII-B Other considerations
VII-B1 Generalizing C3 Heuristics
We propose in this work heuristics for both schedule prioritization (SP) and resource partitioning (RP) for two concurrent kernels. Our SP heuristic can be extended to more kernels, that is, instead of two kernels, runtime can prioritize scheduling of multiple kernels again in the order of their resource requirements (number of workgroups), low to high. Similarly, our RP heuristic timing analysis can be extended to more kernels. As for two kernels, we can slowdown multiple kernels using the proposed analytical model and assess CU allocation that leads to best performance. However, this model does not factor in the increasing memory interference with more concurrent kernels and further investigation can be necessary in this regard. We leave evaluating these heuristics for more concurrent kernels to future work.
VII-B2 Caching Considerations
In realistic applications, usage of DMA offloads can have interesting caching implications. As an example, on MI300X, a DMA transfer will not observe dirty data from a preceding computation which leaves dirty data in L2. In such cases, a cache flush from L2 to AMD Infinity Cache can be necessary before invoking a DMA transfer. Alternately, preceding kernels can proactively write to common cache levels between compute and DMA engines hiding most of this cost.
VII-B3 Inter-Node Communication Considerations
Our study primarily addresses intra-node communication within an MI300X Infinity Platform node. While this is most prominent for inference and small-scale training, large-scale distributed ML training involves inter-node communication as well. In such cases, ML algorithms attempt to use both intra and inter node communication in tandem with attempting to maximize former as much as possible for better performance. In such cases, even with a focus on intra-node communication only, ConCCL can be useful end-to-end.
VII-B4 Global/Local Optimization
Our proposed heuristics/techniques can in some sense be classified as local optimization strategies. That is, we identify given two or more concurrent kernels, which of schedule prioritization or resource partitioning gets better performance. We understand that for better end-to-end performance for an application, with many kernels, such local decisions have to be combined at global-level to get better performance. Prior works in this regard [20] can be used in tandem with our work to do so.
VII-B5 Power Considerations
As GPUs get more capable, power constraints are increasingly at the forefront in their design. A power-agnostic scheduler could, by over-employing C3, lower performance by causing GPU power to be stressed leading to power management events. Alternately, a more power-aware scheduler can employ C3 more judiciously by prioritizing concurrency for complementary power kernels. We leave design of such heuristics to future work.
VIII Related Work
Given the prevalence of C3 in ML, optimized C3 is key to continued scaling of ML. In our work, we methodically study C3, analyze resulting interference and most importantly, without making any invasive changes in existing system software and hardware, we identify techniques to use available software and hardware resources in optimized way to maximize benefit of C3. Computation-Communication Kernel Overlap: Several GPU communication kernel libraries have been developed which improve collective performance in isolation - RCCL [14], MSCCLang [21], and more [22, 23, 24]. However, none of these take any special measures to mitigate any interference where collective kernels execute concurrently with computation kernels which are abundant in ML executions. Our work instead systematically studies execution of RCCL [14] collectives (which incorporates optimizations from other libraries for an AMD MI300X GPU) in concurrence with compute operations and takes necessary measures to mitigate interference when possible. While some prior works study this overlap algorithmically, they do not provide insights into the resource contention therein [11]. In this work, we not only study these contentions, but also develop two optimizations including, schedule prioritization and compute (CU) partitioning, that improve end-to-end performance of C3. While there have been numerous works on improving concurrent kernel execution on GPUs via appropriate resource partitioning [25, 26], prioritization [20] and concurrency-aware kernels [27, 28, 29], they primarily focus on multiple compute kernels. Our work applies these techniques to coarse-grain compute-communication kernels. NanoFlow [3] also performs resource partitioning for C3; however, they create very fine-grained kernels by slicing ML input from micro into nano-batches, and find necessary resource allocation from a huge and complex search space unlike our simple yet useful lookup based heuristic.
Communication Offload: Works which improve C3 performance usually offload communication to dedicated customized accelerators on GPUs, for example, ACE [30] or require extensive hardware support such as compute-capable switches [31]. We consider offload but by leveraging existing data-movers (DMAs), requiring no hardware changes. While dedicated accelerators/compute-enhanced switches can theoretically have lesser memory contention via internal buffering, in practice this is challenging due to buffer overflows which can only be fixed with large buffers that are impractical for large-sized collectives.
Communication Offload to DMAs: There are also work which offload communication to DMA engines. For example, MSCCL++ [24] offloads larger size collective operations to DMA engines, however it initiates DMA call from GPU involving CU (through proxy channel in CPU), which in our work we prefer to reserve only for compute operations. ARK [5] suggests communication offload to DMA but emphasizes on DMA call overhead time and involves GPU threads to reduce such overhead either by significant SW changes or by addition of new DMA prototype; we, in contrast rely on existing software stack and HW and still show benefit of sDMA offload. Some of the works [32] leverage DMAs for fine-grained C3. However, unlike this work, they provide limited insights into challenges with such overlap and therefore provide limited insights into the potential for DMAs alone to help overcome C3 interference like this work highlights with ConCCL. Other works require hardware modifications to initiate DMA transfers in a fine-grained manner [19].
Fine-grained Computation-Communication Overlap: Other interesting works discuss techniques for fine-grain overlap of compute/communication to improve concurrency [33, 34, 35, 3, 19, 32]. While such fine-graining techniques are promising, they necessitate pervasive changes in SW/HW to realize overall performance benefit. For example, Fused Computation-Collective [33] fuses communication operation directly in GEMM kernel requiring exponential implementations in GEMM library.
IX Conclusion
In this work, we carefully characterize the performance of concurrent computation and communication (C3), an important ML paradigm, on state-of-art MI300X GPU and observe that speedups attained are, not unexpectedly, on average only 21% of ideal speedup possible due to interference between concurrent GPU kernels. To improve performance of this important primitive, we first evaluate dual strategies of schedule prioritization and careful resource partitioning of compute units on GPUs to push performance attained with C3 (on average 42% of ideal speedup). Additionally, we provide heuristics that can guide a GPU runtime to harness these strategies. To further improve speedups for C3, we offload communication to DMA engines via building ConCCL (Concurrent Communication Collectives) proof-of-concepts. We demonstrate how ConCCL closes the gap between realized and ideal speedup for C3 (on average about 72% of ideal speedup is realized, up to 1.67 speedup). We believe DMA engines have an important role to play for C3 performance and argue for their continued betterment.
Acknowledgment
AMD, the AMD Arrow logo, AMD CDNA, AMD Instinct, AMD ROCm, AMD Infinity Cache, AMD Infinity Fabric, and combinations thereof are trademarks of Advanced Micro Devices, Inc. Other product names used in this publication are for identification purposes only and may be trademarks of their respective companies.